40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Feb 3 15:57:52 2014, Steve, Update, Green Locking, ETMX green power ETMXgreen.jpgXgreenOven36C.jpg
    Reply  Tue Feb 4 01:02:03 2014, Koji, Update, Green Locking, ETMX green power Screenshot-Untitled_Window.png
    Reply  Tue Jan 19 17:27:58 2016, gautam, Update, Green Locking, AUX X power investigations X_END_POWER_BUDGET.JPG
Message ID: 9592     Entry time: Mon Feb 3 15:57:52 2014     Reply to this: 9595   11936
Author: Steve 
Type: Update 
Category: Green Locking 
Subject: ETMX green power 

There was 0.2 mW green at the X end.

The doubling oven temp was changed from 37.5 to 36 degrees C

Power at green shutter 3 mW  The alignment was not touched.

Attachment 1: ETMXgreen.jpg  308 kB  | Hide | Hide all
ETMXgreen.jpg
Attachment 2: XgreenOven36C.jpg  147 kB  | Hide | Hide all
XgreenOven36C.jpg
ELOG V3.1.3-